the sides of a triangle are (x+2), (3x+1) and (x+7) units. find the units of x if the perimeter of the triangle is 140 units

Answer by ikleyn(52781)

.
the sides of a triangle are (x+2), (3x+1) and (x+7) units. find the units of x if the perimeter of the triangle is 140 units
~~~~~~~~~~~~~~~


Your equation is


    (x+2) + (3x+1) + (x+7) = 140.


Simplify and solve for x.


    5x = 140 - 2 - 1 - 7 = 130.

     x = 130/5 = 26.


The sides are  (26+2) = 28;  3*26+1 = 79;  26+7 = 33.

Since   79 > 28 + 33,  the thriangle inequality is   V I O L A T E D,  which means that

        SUCH A TRIANGLE DOES NOT EXIST.


Your problem is a   F A K E.